# Fluctuation Theorem and Optimal Control of an Active Tracking Particle with Information Processing

### v3

@@ Abstract
Suppose that during each time step, $t_{k}\rightarrow t_{k+1}$, the energy consumed for sustaining the particle activity is $E_{A}^{0}$, which is usually proportional to $v_{0}^{2}$ with $v_{0}$ as self-propulsion speed, the energy of applied magnetic field is $E_{B}=E_{B}^{0}\cdot(B/B_{0})^{2}=E_{B}^{0}\hat{B}^{2}$, and the energy for measurement is $E_{M}=-E_{M}^{0}\ln\epsilon$, which increases with the decreasing error. Then the ensemble-averaged total energy consumed for arriving at the destination is,
$\langle\mathcal{E}\rangle$ $=$ $\frac{\mathcal{D}E_{A}^{0}}{2p(n^{\prime}_{k}=R)-1}~\left[1+\mathcal{B}\hat{B}^{2}\cdot p(m_{k}=L)-\mathcal{M}\ln\epsilon\right],$
where $p(m_{k}=L)=1/2$ denotes the probability of the measured particle orientation as $m_{k}=L$, and the referenced magnetic energy, $\mathcal{B}=E_{B}^{0}/E_{A}^{0}$, and referenced measurement energy, $\mathcal{M}=E_{M}^{0}/E_{A}^{0}$, are two system parameters.
@@
Note that the measurement energy, $E_{M}$, increases with the decreasing error in measurements, and meanwhile, the decreasing error can shorten the total time through biasing the probability $p(n^{\prime}_{k}=R)$ and thus lowers the energy spent on the active motion and the control field; such a competition can lead to the issue of (information) robustness - (measurement) energy tradeoff, commonly discussed in biological systems [59, 60, 61]. As shown in Fig. 3(a), for system parameters of $\mathcal{B}=1/2$ and $\mathcal{M}=1$, the total consumed energy can have a local minimum (also the global minimum in this case) at $(\epsilon^{\mathrm{opt}},\hat{B}^{\mathrm{opt}})$, which serves as the optimal control variables for minimizing the total energy consumption. By increasing the referenced magnetic energy, e.g., $\mathcal{M}=30$ in Fig. 3(b), the local minimum is no longer the global one, where the global minimum is now located at $\epsilon=1$ which is a natural upper bound of the measurement error. Note that the global minimum located at $\epsilon=1$ is not necessarily an extremum (usually not). Meanwhile, the situation can become more complicated at large $\mathcal{B}$, where there can be two local minima, as shown in Fig. 3(c). Then the global minimum is chosen among the local minima and the one at $\epsilon=1$, which is shown in Fig. 3(d,e). In this case, the system can exhibit a transition of controllable variables $(\epsilon^{\mathrm{opt}},\hat{B}^{\mathrm{opt}})$ depending on the system parameters $(\mathcal{B},\mathcal{M})$, where the guiding rule is to minimize the total energy consumption. There are two important limits to discuss about, $\mathcal{M}=0$ and $\mathcal{M}\rightarrow\infty$. For the measurement parameter, $\mathcal{M}=0$, the measurements of optimal controls are error-free, i.e., $\epsilon^{\mathrm{opt}}=0$, and we can obtain the optimal external field follows the relation, $\hat{B}^{\mathrm{opt}}_{\epsilon=0}\simeq\frac{1}{2}W(4/\mathcal{B})$ [$W(x)$ as the principal branch of Lambert W function] at $\mathcal{B}\rightarrow 0$ [blue line in Fig. 3(f)], and $\hat{B}^{\mathrm{opt}}_{\epsilon=0}\simeq\sqrt{2/\mathcal{B}}$ at $\mathcal{B}\rightarrow\infty$ [red line in Fig. 3(f)]. For the measurement parameter, $\mathcal{M}\rightarrow\infty$, the measurements of optimal controls are wrong, i.e., $\epsilon^{\mathrm{opt}}=1$, and we can obtain the optimal external field follows the relation, $\hat{B}^{\mathrm{opt}}_{\epsilon=1}\simeq\frac{1}{2}W(8/[\mathcal{B}\mathcal{(}1-u_{p})])$ at $\mathcal{B}\rightarrow 0$ [blue line in Fig. 3(g)], and $\hat{B}^{\mathrm{opt}}_{\epsilon=1}\simeq\sqrt{2/\mathcal{B}}$ at $\mathcal{B}\rightarrow\infty$ [red line in Fig. 3(g)].
@@
In this work, we develop a smart active model where activity, thermal fluctuation and information processing are incorporated, in analogy of a run-and-tumble bacterium equipped with a ‘brain’. After deriving the generalised fluctuation theorem setting the relation between the entropy production and the information flow in the system, we discuss the optimal feedback control strategies regarding the first passage steps and the total energy consumption of the system. The performed analyses can be directly generalised to other systems with different optimisation goals such as minimal excess work [62, 63] and dissipated heat [63, 64]. We anticipate the concepts developed here to be applied for uncovering the physical rules underlying more complicated responses of regulatory living systems including collective dynamics [65, 66, 67, 68, 69, 70, 71], and also to be useful in future industrial designs of smart active systems with desired physical responses.
@@ Acknowledgements.
F.M. acknowledges the supports by the National Natural Science Foundation of China (Grant No. 12275332 and 12447101), Wenzhou Institute (Grant No. WIUCASQD2023009), and Beijing National Laboratory for Condensed Matter Physics (Grant No. 2023BNLCMPKF005).
     Appendix
The probability, $p(n^{\prime}_{k})$. In the table, $p(n^{\prime}_{k})$, utilised for analysing the mean first passage steps and the mean total energy consumption, depends on both the magnetic field and the measurement error explicitly. In the table, $p_{\mathrm{eq}}^{0}(n_{k}=L,R)=p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L,R)=1/2$ denotes the equilibrium distribution of $n_{k}$ and $n_{k}^{\prime}$ without applying magnetic field, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$, respectively, without applying magnetic field. $p_{\mathrm{eq}}^{B}(n_{k}^{\prime}=L)=\exp[-mB/(k_{B}T)]/\{\exp[-mB/(k_{B}T)]+\exp[mB/(k_{B}T)]\}$ and $p_{\mathrm{eq}}^{B}(n_{k}^{\prime}=R)=1-p_{\mathrm{eq}}^{B}(n_{k}^{\prime}=L)$ denote the equilibrium distribution of $n_{k}^{\prime}$ with applied magnetic field, and the transition rate is obtained as in the Kramers process [72, 73], $k_{M}=k_{M}^{L\rightarrow R}+k_{M}^{R\rightarrow L}=k_{0}\cosh[mB/(k_{B}T)]$ with applied magnetic field.
$n^{\prime}_{k}$ $p(n^{\prime}_{k})$
$L$ $p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)+p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)\epsilon$
$p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)-p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)(1-\epsilon)+[p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)+p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)(1-\epsilon)$
$p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)\epsilon$
$R$ $p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)-p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)\epsilon$
$p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=R)+p_{\mathrm{eq}}^{0}(n^{\prime}_{k}=L)\cdot e^{-k_{0}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)(1-\epsilon)+[p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=L)(1-\epsilon)$
$p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)+p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)\cdot e^{-k_{M}\Delta t}]p_{\mathrm{eq}}^{0}(n_{k}=R)\epsilon$
Table 1: The probability, $p(n^{\prime}_{k})$.
The probability, $p(n^{\prime}_{k}|n_{k},m_{k})$. Here we show an important probability utilised in the main text, $p(n^{\prime}_{k}|n_{k},m_{k})$. For other probabilities, one can refer to supplemental materials for their explicit expressions. [53].
$n_{k},m_{k},n^{\prime}_{k}$ $p(n^{\prime}_{k}|n_{k},m_{k})$
$L,R,L$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot e^{-k_{0}\Delta t}$
$L,R,R$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot e^{-k_{0}\Delta t}$
$R,R,L$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)]\cdot e^{-k_{0}\Delta t}$
$R,R,R$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)]\cdot e^{-k_{0}\Delta t}$
$L,L,L$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)+[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)]\cdot e^{-k_{M}\Delta t}$
$L,L,R$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)-[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)]\cdot e^{-k_{M}\Delta t}$
$R,L,L$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=L)-[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)]\cdot e^{-k_{M}\Delta t}$
$R,L,R$ $p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)+[1-p_{\mathrm{eq}}^{B}(n^{\prime}_{k}=R)]\cdot e^{-k_{M}\Delta t}$
Table 2: The probability, $p(n^{\prime}_{k}|n_{k},m_{k})$ in the regulatory process.

### v4

@@ Abstract
Suppose that during each time step, $t_{k}\rightarrow t_{k+1}$, the energy consumed for sustaining the particle activity is $E_{A}^{0}$, which is usually proportional to $v_{0}^{2}$ with $v_{0}$ as self-propulsion speed, the energy of applied magnetic field is $E_{B}=E_{B}^{0}\cdot(B/B_{0})^{2}=E_{B}^{0}\hat{B}^{2}$, and the energy for measurement is $E_{M}=E_{M}^{0}\langle I(m_{k}:n_{k})\rangle=E_{M}^{0}[\ln 2+\epsilon\!~\mathrm{ln}\!~\epsilon+(1-\epsilon)\ln(1-\epsilon)]$. Note that it is different from the case of position measurement; in that case, a particle is located at $x$ in a 1D confinement of size $L$ and its position is measured with error $\Delta x$, so the relative error is $\epsilon=\Delta x/L$ and the measurement energy is constructed simply as $E_{M}=-E_{M}^{0}\ln\epsilon$ [49, 50]. Then the ensemble-averaged total energy consumed for arriving at the destination is,
$\langle\mathcal{E}\rangle$ $=$ $\frac{\mathcal{D}E_{A}^{0}}{2p(n^{\prime}_{k}=R)-1}~\left[1+\frac{1}{2}\mathcal{B}\hat{B}^{2}+\mathcal{M}\langle I(m_{k}:n_{k})\rangle\right],$
where the referenced magnetic energy, $\mathcal{B}=E_{B}^{0}/E_{A}^{0}$, and referenced measurement energy, $\mathcal{M}=E_{M}^{0}/E_{A}^{0}$, are two system parameters. Although the measurement energy is symmetric with respect to $\epsilon=1/2$, i.e., $E_{M}(\epsilon)=E_{M}(1-\epsilon)$, the symmetry gets broken when taking into account of the other two energy contributions, and minimisation of $\langle\mathcal{E}\rangle$ would always give $\epsilon\leq 1/2$.
@@
Note that the measurement energy, $E_{M}$, increases with the decreasing error for $0\leq\epsilon\leq 1/2$, and meanwhile, the decreasing error can shorten the total time through biasing the probability $p(n^{\prime}_{k}=R)$ and thus lowers the energy spent on the active motion and the control field; such a competition can lead to the issue of (information) robustness - (measurement) energy tradeoff, commonly discussed in biological systems [66, 67, 68]. As shown in Fig. 3(a), for system parameters of $\mathcal{B}=1/2$ and $\mathcal{M}=1$, the total consumed energy can have a local minimum (also the global minimum in this case) at $(\epsilon^{\mathrm{opt}},\hat{B}^{\mathrm{opt}})$ with a very small measurement error, which serves as the optimal control variables for minimising the total energy consumption. By increasing the measurement parameter, e.g., $\mathcal{M}=30$ in Fig. 3(b), there is still one local minimum serving as the global minimum, but with a relatively large measurement error, $\epsilon=0.45$. Meanwhile, the situation can become more complicated at large $\mathcal{B}$, where there can be two local minima, as shown in Fig. 3(c) (see supplemental materials for more about the extrema including saddle points [51]). Then the global minimum is chosen between the two local minima, which is shown in Fig. 3(d,e). In this case, the system can exhibit a transition of controllable variables $(\epsilon^{\mathrm{opt}},\hat{B}^{\mathrm{opt}})$ depending on the system parameters $(\mathcal{B},\mathcal{M})$, where the guiding rule is to minimise the total energy consumption. Note that the optimal field $\hat{B}^{\mathrm{opt}}$ appears largely insensitive to the measurement parameter $\mathcal{M}$, except near the strategy bifurcation points, while $\epsilon^{\mathrm{opt}}$ is shown sensitive to $\mathcal{M}$, and such difference roots in the different energy forms between the magnetic energy $E_{B}=E_{B}^{0}\hat{B}^{2}$ (quadratic dependence on $\hat{B}$) and the measurement energy $E_{M}=E_{M}^{0}[\ln 2+\epsilon\!~\mathrm{ln}\!~\epsilon+(1-\epsilon)\ln(1-\epsilon)]$ (logarithmic dependence on $\epsilon$). There are two important limits to discuss about, $\mathcal{M}=0$ [Fig. 3(f)] and $\mathcal{M}\rightarrow\infty$ [ Fig. 3(g)]. For the measurement parameter, $\mathcal{M}=0$, the measurements of optimal controls are error-free, i.e., $\epsilon^{\mathrm{opt}}=0$, and we can obtain the optimal external field follows the relation, $\hat{B}^{\mathrm{opt}}_{\epsilon=0}\simeq\frac{1}{2}W(4/\mathcal{B})$ [$W(x)$ as the principal branch of Lambert W function] at $\mathcal{B}\rightarrow 0$, and $\hat{B}^{\mathrm{opt}}_{\epsilon=0}\simeq\sqrt{2/\mathcal{B}}$ at $\mathcal{B}\rightarrow\infty$. For the measurement parameter, $\mathcal{M}\rightarrow\infty$, the measurements are useless (or not performed), i.e., $\epsilon^{\mathrm{opt}}=1/2$, and we can obtain the optimal external field follows the relation, $\hat{B}^{\mathrm{opt}}_{\epsilon=1}\simeq\frac{1}{2}W(8/\mathcal{B})$ at $\mathcal{B}\rightarrow 0$, and $\hat{B}^{\mathrm{opt}}_{\epsilon=1}\simeq\sqrt{2/\mathcal{B}}$ at $\mathcal{B}\rightarrow\infty$.
Moreover, the system can be generalised to become other important information-regulated active matter systems, e.g., active information engine [69, 70]. If the active particle itself is regarded as a cargo, then the system setup can be applied to describe the cargo transportation process. By introducing $\mathcal{W}=\mathcal{D}\cdot\zeta v_{0}\Delta x$ as the effective work for transporting the particle from location $x=0$ to destination $x=\mathcal{D}\Delta x$, where $\zeta$ denotes the friction coefficient of the particle and $\zeta v_{0}$ denotes the effective driving force on a passive loading, we can define the work efficiency as $\eta=\mathcal{W}/\langle\mathcal{E}\rangle$, with $\langle\mathcal{E}\rangle$ the averaged total energy consumption introduced above. In this simple setup of cargo transportation, the total work is constant, $\mathcal{W}=\mathcal{D}\cdot\zeta v_{0}\Delta x$, so the goal of ‘maximising work efficiency’ is equivalent to the goal of ‘minimising total energy consumption’, thus, the performed analyses of the minimising the total energy consumption above can be directly utilised for discussing about the work efficiency. One can also generalise the setup to deal with more complicated information engines in the future, e.g., adding passive or actively-competing loadings on the active tracking particle and discuss about the work efficiency in different setups.
@@
In this work, we develop a smart active model where activity, thermal fluctuation and information processing are incorporated, in analogy of a run-and-tumble bacterium equipped with a ‘brain’. After discussing about the entropy production and the information flow of the system, we illustrate different optimal feedback control strategies regarding the first passage steps and the total energy consumption of the system, with a simple generalisation to active information engine. Specifically, by setting the optimisation goal as minimising the total energy consumption, there are interesting findings, including the (information) robustness - (measurement) energy tradeoff and the strategy transition, due to the interplay between active motion, measurements and feedback controls. Such energetics analyses can be generalised for investigating physical responses of other ‘smart’ active systems, where the information-based regulations can be quantified to provide intuitive understandings, where one may find relevant tradeoffs of competing effects and similar strategy transitions. Possible future connections with experiments on controllable active tracking particles are promising, e.g., platinum-nickel-gold nanorods [54], gold-polystyrene Janus particles [55] and metal-dielectric Janus particles [56, 57], however, necessary data including measurement energy, fuel consumption, control energy, etc., needs to be measured for comparisons. Moreover, we also anticipate this study to inspire more theoretical works on formulating complex phenomena of ‘smart’ systems, e.g., collective dynamics of information-processed systems [71, 72, 73, 74, 75], and also to be useful in future industrial designs of smart active systems with desired physical responses.
@@ Acknowledgements.
F.M. acknowledges the supports by the National Natural Science Foundation of China (Grant No. 12275332 and 12447101), Wenzhou Institute (Grant No. WIUCASQD2023009), UCAS Xiaomi Youth Fellowship, and Beijing National Laboratory for Condensed Matter Physics (Grant No. 2023BNLCMPKF005).